OpenStudy (anonymous):

A $6000 investment earns $690 simple interest in two years. What is the annual interest rate? A. 1.15% B. 5.75% C. 11.5% D. 57.5%

jimthompson5910 (jim_thompson5910):

I = P*r*t 690 = 6000*r*2 solve for r and convert to a percentage